# Neutron star evolution by combining discontinuous Galerkin and finite volume methods

### v1

@@ IV.5.3 DG+FV/FD (U) simulation:
Last, we compare the TEI perfromance for the DG+FV/FD simulations with different levels of h-refinement in Fig. 17. This provides another crucial test, as with increasing h-refinement the TEI should hone in on the troubled regions near the star surface, thus using FV/FD in a smaller region. We see this successfully carried out in Fig. 17, where the lower h-refined simulation in the left plots show a thicker region of FV/FD elements, as compared to the counterparts from the higher h-refined simulation on the right. The top plots correspond to the time when the star has expanded fully for the first time, and $\rho_{0,\mathrm{C}}$ has reached its first minimum, while the bottom plots are for the following maximum in $\rho_{0,\mathrm{C}}$, when matter has fallen in. For both these states of the star, the times for the two resolutions of $l=3$ and $l=4$ has been chosen while accounting for the time offset in the dynamics of the two resolutions to achieve a faithful comparison. So, we can conclude from these plots that the TEI perfroms properly and retains DG in more elements as the element size decreases with increasing h-refinement.
@@ IV.6 Boosted configuration (B)
Figure 19: Portion of the mesh in the $xy$-plane from the (B) simulation with $v_{\mathrm{CoM}}=0.5$ using the DG+FV/FD evolution method at four different times. Top left: $t=0M_{\odot}$; top right: $t=40M_{\odot}$; bottom left: $t=88M_{\odot}$; and bottom right $t=132M_{\odot}$. The elements with denser, equally spaced grid points, are the ones using FV/FD. The DG elements are the ones with sparser, unequally spaced grid points. The colormap shows the rest mass density $\rho_{0}$. We see that the TEI nicely track the surface of the star and use FV/FD only in that region. Furthermore, we also see the adaptive mesh refinement spheres track the star. The simulation continues smoothly even beyond the central cubical region.
[…]
@@
We mainly concentrate here on the efficiency of our TEI. In the wireframe colormap plot of $\rho_{0}$ in the $xy$-plane for $v_{\mathrm{CoM}}=0.5$ in Fig. 19, the left top plot is the initial setup. Here, we have again started out by setting only elements near the star surface to evolve with FV/FD as in Sec. IV.3.2. We are a bit more cautious than for (S) with the initial FV/FD setup, as we know that (B) will start moving, unlike (S). Hence, not only do we set elements that contain the star surface to FV/FD, but also their the nearest (von Neuman) neighbors. These results in a comparatively thicker initial band of FV/FD elements at the surface of the star. The star is placed at a corner of the $xy$-plane and the $\vec{v}_{\mathrm{CoM}}$ is imparted diagonally along the plane to keep the star for longer in the cubical patch. The TEI kicks in during the evolution, and we see that only the surface of the star is being evolved with FV/FD in the next plot at $t=40M_{\odot}$ on the top right. This switch to FV/FD only on the surface happens early, near the very beginning of the simulation. There is a small amount of matter with density $10^{-3}\rho_{0,\mathrm{C}}(t=0)$ that has spread out, which is also evolved with FV/FD. We also see that the refinement spheres follow the star center. In the lower left plot at time $t=88M_{\odot}$, we see that the star has reached the corner of the central cubical region, which is the best resolved region. The TEI continues detecting the troubled region and evolve this with FV/FD scheme, while maintaining DG in the rest of the mesh. In the last (lower right) plot at $t=132M_{\odot}$, we see that the star has crossed over from the central cube to the adjacent cubed sphere region. Since these regions are less resolved, and since interpolation takes place at the surface where the central cube and the adjacent cubed spheres meet, there is a loss of accuracy when matter flows from the central cube to the adjacent cubed sphere patches. Hence, we do not plan to infer any significant conclusions from the results obtained after the point when the majority of the star has moved out of the central cube. Having said that, we observe that Nmesh is able to gracefully handle the transition of the star into these lower resolved region, and the simulation keeps running smoothly, with the TEI still maintaining the FV/FD scheme only in the narrow band near the star surface.
@@
Lastly, in Fig. 20 we look at the central density $\rho_{0,\mathrm{C}}$(left) and the L2 norm of the Hamiltonian constraint in the central cube for the $v_{\mathrm{CoM}}=0.1$ case for the duration of time for which the star remains in the central cube. We choose $v_{\mathrm{CoM}}=0.1$ so that the star remains in the central cube for much longer, and hence provides reliable results for a longer duration. We plot results for the two resolutions with central refinement sphere h-refinment at $l=4$ and $l=5$. We see both the $\rho_{0,\mathrm{C}}$ and Hamiltonian constraint behavior improves with resolution. The results however are affected by the fact that the $l=4$ resolution is too low and hence the entire star is evolved with FV/FD only, as the elements are too large and all contain troubled regions. We leave exploration of behavior with higher resolutions to future work.
@@ V Conclusions
In this work we present a new FV/FD method that can be easily combined with a DG scheme. The aim is to construct a hybrid scheme which uses the DG method in all regions with smooth matter fields, but which switches to FV/FD in regions with non-smooth matter, such as shocks or star surfaces.
@@
The strength of our FV/FD method is that it parallelizes as well as the DG method, because we exchange information between neighboring elements only from the surface of the element. Through this, we minimize the amount of data that needs to be exchanged, and we also ensure that information needs to be exchanged only between nearest neighbors. This should allow for high scalability when parallelizing, even while using the FV/FD method. This is a key difference between our method and other such hybrid schemes. For example, in [114] or [87, 65, 97], while the hybridization concept is similar, the FD elements still need ghostzones consisting of points deeper than just the surface from the adjacent elements to exchange information. Futhermore, as soon as one uses a mesh that contains more than simple cuboid patches, the FD method in [114] or [97] requires information exchange between all 26 neighbors, while our method requires information exchange only between the six nearest neighbors.
@@
The price for not having additional ghost points is that near element boundaries our FV/FD method is only second order accurate. While second order convergence may not sound too impressive, we note that in shocked regions we cannot expect more than first order convergence in any case [92]. Thus, our method is adequate when used in elements with shocks. In this way we obtain benefits of both FV/FD and DG as suggested in [114, 87, 65].
@@
We run a battery of tests with our Nmesh program to verify how well our new method works in paractice. We find that the FV/FD method indeed reaches second order accuracy for tests with smooth scalar fields, and that it also performs well in a shock tube test. We then move on to tests with neutron stars in full GRHD. As already described in [76], neutron stars are more challenging due to the presence of vacuum or low density regions outside the star. We describe in detail the positivity limiters and the atmosphere treatment that we have develpoded to deal with this challenge. Our combination of limiters and atmosphere are successful in the sense that they enable us to evolve neutron stars with the DG method alone. However, the stars are then plagued by Gibbs oscillations [88, 89, 90, 91] that cause the stars to pulsate. This is especially prominent when the stars are moving. We therefore introduce a troubled-element indicator (TEI) that uses a list of trouble criteria designed to detect elements that are not giving good results with the DG method. Using this TEI, we build a hybrid scheme that switches between DG and FV/FD as needed. In particular, if an element is marked as troubled and was using DG, we redo the time step with FV/FD. In this way we try to avoid Gibbs oscillations from ever occuring. In case a troubled element is using FV/FD already we remain with FV/FD. On the other hand, if an element is marked as not troubled for eleven consecutive steps and was using FV/FD, we switch to DG, while we remain with DG in case the element was already using it. In DG elements we use only half the number of points per direction as in FV/FD elements, because DG is able to provide higher accuracy results at lower cost due to its higher rate of convergence.
@@
We believe that this is the first time a DG method has been used to evolve the more challenging boosted star (B) and migration (U) cases in full 3D. In case (U), an overdense unstable star gets perturbed to rapidly expand for a while, but later recontracts. The ensuing strong star pulsations contain actual shocks, besides the non-smooth star surface. Again our hybrid method seems capable of selecting the correct method. The bulk of the star interior and regions further from the star still use the DG method as desired, while FV/FD is selected near the surface and in shocked regions.
@@ Acknowledgements.
It is a pleasure to thank Nils Deppe for helpful discussions. This work was supported by NSF grants PHY-2136036 and PHY-2408903. We also acknowledge usage of computer time on the HPC clusters KoKo and Athene at Florida Atlantic University, as well as on Bridges-2, Expanse and Anvil under ACCESS allocations PHY220018 and PHY230109. The figures in this article were created using matplotlib [129, 130] and TikZ [131].
@@ Appendix A Important parameter specifications
Here we state the values of the most important parameters for the different classes of simulations for (S), (P), (U) and (B) that we have discussed in Sec. IV. These are listed in Table 3. By classes, we mean the different sections described above. Hence, they are simply refered to by their section numbers as: §IV.3.1: (S) with pure DG; §IV.3.2: (S) with DG+FV/FD, §IV.4: (P) with pure DG; §IV.5.1: (U) with pure FV/FD and varying p-refinement; §IV.5.2: (U) with pure FV/FD and varying h-refinement; §IV.5.3: (U) with DG+FV/FD; §IV.6: (B) with DG+FV/FD. We are using the shorthand notation “§” in place of “Sec.” here to denote the sections. Of course Nmesh  has many more parameters that we choose not to list here. When same parameter specification is used across multiple classes, we group those classes together in the table. In cases same specification is used for all simulations, we simply say “all” in the “classes” column. Some parameter choices are used for some runs, which are not used for others entirely. In such cases, it is likely that some other parameter is playing an analogous role. For example, when we set the parameter GHG_gammas_type to constants, then we are using the constant constraint damping, and we specify the values of the parameters GHG_gamma0 $=0.1$, GHG_gamma1 $=-1$ and GHG_gamma2 $=1$. However, when we set GHG_gammas_type to 3gaussians, as in we set the position dependent triple Gaußian constraint damping functions from Eq. (5), then we have to use the counterparts parameters for these functions. We set these values to:
@@
$A^{(0)}_{(2)}$ and $A^{(2)}_{(2)}$ are set to 00 in our case, as unlike in [97], we have only one star in our simulations. We name the parameters corresponding to the constants $C_{(0)}$, $A^{(0)}_{(1)}$ and $w^{(0)}_{(1)}$ as GHG_3gauss_gamma0_C, GHG_3gauss_gamma0_A1 and GHG_3gauss_gamma0_w1 respectively, and the name of the rest of the parameters follow suite. Owing to the large number of constants, we do not repeat these position dependent constarint damping parameters in the table. All the run classes that sets GHG_gammas_type = 3gaussians use these values for the constants.
@@
by setting amplitude of damped_wave to 00.
@@
GRHD_atmo_cut: Factor $f_{\mathrm{atmo,cut}}$ $\mathtt{1.01\times 10^{-12}}$IV.3.1, §IV.4,
from Sec. III.5.1. $\mathtt{10^{-9}}$IV.5.1, §IV.3.2, §IV.5.2, §IV.5.3, §IV.6
@@
GRHD_atmo_level: Factor $f_{\mathrm{atmo,level}}$ $\mathtt{10^{-12}}$IV.3.1, §IV.4,
from Sec. III.5.1. $\mathtt{10^{-11}}$IV.5.1, §IV.3.2, §IV.5.2, §IV.5.3, §IV.6
@@
GRHD_troublescore_cut: Factor $f_{\mathrm{cut,low}}$ related to Eq. (50). $\mathtt{10^{-4}}$IV.3.2, §IV.5.3, §IV.6

### v2

@@ IV.5.3 DG+FV/FD (U) simulation
Last, we compare the TEI performance for the DG+FV/FD simulations with different levels of h-refinement in Fig. 17. This provides another crucial test, as with increasing h-refinement the TEI should hone in on the troubled regions near the star surface, thus using FV/FD in a smaller region. We see this successfully carried out in Fig. 17, where the lower h-refined simulation in the left plots show a thicker region of FV/FD elements, as compared to the counterparts from the higher h-refined simulation on the right. The top plots correspond to the time when the star has expanded fully for the first time, and $\rho_{0,\mathrm{C}}$ has reached its first minimum, while the bottom plots are for the following maximum in $\rho_{0,\mathrm{C}}$, when matter has fallen in. For both these states of the star, the times for the two resolutions of $l=3$ and $l=4$ have been chosen while accounting for the time offset in the dynamics of the two resolutions to achieve a faithful comparison. So, we can conclude from these plots that the TEI performs properly and retains DG in more elements as the element size decreases with increasing h-refinement.
@@ IV.6 Boosted configuration (B)
Figure 19: A portion of the mesh in the $xy$-plane from the (B) simulation with $v_{\mathrm{CoM}}=0.5$ using the DG+FV/FD evolution method at four different times. Top left: $t=0M_{\odot}$; top right: $t=40M_{\odot}$; bottom left: $t=88M_{\odot}$; and bottom right $t=132M_{\odot}$. The elements with denser, equally spaced grid points, are the ones using FV/FD. The DG elements are the ones with sparser, unequally spaced grid points. The colormap shows the rest mass density $\rho_{0}$. We see that the TEI nicely tracks the surface of the star and uses FV/FD only in that region. Furthermore, we also see the adaptive mesh refinement spheres track the star. The simulation continues smoothly even beyond the central cubical region.
The last case we study is the boosted star (B), that has an initial boost applied to (S). We evolve two (B) setups with boost applied such that the CoM of the star has a velocity of $v_{\mathrm{CoM}}=0.1$ and $0.5$ diagonally along the $xy$-plane. (B) again proves too difficult to evolve with pure DG, as large unphysical Gibbs oscillations develop that reduce the accuracy significantly. Hence, we again turn to the DG+FV/FD method. Almost all aspects of the simulation setups for (B) remain the same as the (S) simulation in Sec. IV.3.2, with a few modifications, as we impart a boost to that very star in the simulations of this section. The larger dimension of $62$ of the cube, compared to the (U) setup and pure DG (S) setup, is chosen for allowing the star to travel longer in the well-refined cubical patch. The resultant initial number of elements both parallel and perpendicular to the $\vec{v}_{\mathrm{CoM}}$ direction are stated in Table 2. This differs for the $0.5$ case for the two directions, but remains the same for $0.1$. The central h-refinement region is refined $l=5$ times for the $v_{\mathrm{CoM}}=0.5$ case and $l=4$ and $l=5$ times for the $v_{\mathrm{CoM}}=0.1$ case we simulate, and the h-refinement falls off successively, the same as in Sec. IV.3.2. This higher level of h-refinement is again set to account for the large dimension of the central box and maintains a comparable number points as the other simulations with DG+FV/FD. Unlike the situations for (S), (P) and (U), where the CoM of the star does not move from its initial position, for (B) we track the star center by following the minimum of the lapse, and use AMR to make the whole h-refinement arrangement follow the moving star. For efficiency, the mesh is not updated after every time step, but rather after regular short intervals of $\delta t=0.7071067811865475$. We also use load balancing and redistribute the elements among all available MPI processes every time the mesh is updated. The balancing is based on the measured time each element spends on doing calculations to determine the workload of each element, with the goal of giving each MPI process the same amount of work. All the other setups remain same as that in Sec. IV.3.2 in all aspects.
@@
We mainly concentrate here on the efficiency of our TEI. In the wireframe colormap plot of $\rho_{0}$ in the $xy$-plane for $v_{\mathrm{CoM}}=0.5$ in Fig. 19, the left top plot is the initial setup. Here, we have again started out by setting only elements near the star surface to evolve with FV/FD as in Sec. IV.3.2. We are a bit more cautious than for (S) with the initial FV/FD setup, as we know that (B) will start moving, unlike (S). Hence, not only do we set elements that contain the star surface to FV/FD, but also their the nearest (von Neuman) neighbors. These result in a comparatively thicker initial band of FV/FD elements at the surface of the star. The star is placed at a corner of the $xy$-plane and the $\vec{v}_{\mathrm{CoM}}$ is imparted diagonally along the plane to keep the star longer in the cubical patch. The TEI kicks in during the evolution, and we see that only the surface of the star is being evolved with FV/FD in the next plot at $t=40M_{\odot}$ on the top right. This switch to FV/FD only on the surface happens early, near the very beginning of the simulation. There is a small amount of matter with density $10^{-3}\rho_{0,\mathrm{C}}(t=0)$ that has spread out, which is also evolved with FV/FD. We also see that the refinement spheres follow the star center. In the lower left plot at time $t=88M_{\odot}$, we see that the star has reached the corner of the central cubical region, which is the best resolved region. The TEI continues detecting the troubled region and evolve this with FV/FD scheme, while maintaining DG in the rest of the mesh. In the last (lower right) plot at $t=132M_{\odot}$, we see that the star has crossed over from the central cube to the adjacent cubed sphere region. Since these regions are less resolved, and since interpolation takes place at the surface where the central cube and the adjacent cubed spheres meet, there is a loss of accuracy when matter flows from the central cube to the adjacent cubed sphere patches. Hence, we do not plan to infer any significant conclusions from the results obtained after the point when the majority of the star has moved out of the central cube. Having said that, we observe that Nmesh is able to gracefully handle the transition of the star into these lower resolved regions, and the simulation keeps running smoothly, with the TEI still maintaining the FV/FD scheme only in the narrow band near the star surface.
@@
Lastly, in Fig. 20 we look at the central density $\rho_{0,\mathrm{C}}$(left) and the L2 norm of the Hamiltonian constraint in the central cube for the $v_{\mathrm{CoM}}=0.1$ case for the duration of time for which the star remains in the central cube. We choose $v_{\mathrm{CoM}}=0.1$ so that the star remains in the central cube for much longer, and hence provides reliable results for a longer duration. We plot results for the two resolutions with central refinement sphere h-refinment at $l=4$ and $l=5$. We see both the $\rho_{0,\mathrm{C}}$ and Hamiltonian constraint behavior improve with resolution. The results however are affected by the fact that the $l=4$ resolution is too low and hence the entire star is evolved with FV/FD only, as the elements are too large and all contain troubled regions. We leave exploration of behavior with higher resolutions to future work.
@@ V Conclusions
In this work we present a new compact FV/FD method that can be easily combined with a DG scheme. The aim is to construct a hybrid scheme which uses the DG method in all regions with smooth matter fields, but which switches to FV/FD in regions with nonsmooth matter, such as shocks or star surfaces.
@@
The strength of our compact FV/FD method is that it parallelizes as well as the DG method, because we exchange information between neighboring elements only from the surface of the element. Through this, we minimize the amount of data that needs to be exchanged, and we also ensure that information needs to be exchanged only between nearest neighbors. This should allow for high scalability when parallelizing, even while using the FV/FD method. This is a key difference between our method and other such hybrid schemes. For example, in [115] or [87, 65, 97], while the hybridization concept is similar, the FD elements still need ghost zones consisting of points deeper than just the surface from the adjacent elements to exchange information. Furthermore, as soon as one uses a mesh that contains more than simple cuboid patches, the FD method in [115] or [97] requires information exchange between all 26 neighbors, while our method requires information exchange only between the six nearest neighbors. In addition, the information exchange between adjacent elements simplifies compared to a traditional FV implementation. Since we always have grid points at the interface, we only need surface interpolation. As a result it is much easier to implement AMR.
@@
The price for not having additional ghost points is that near element boundaries our compact FV/FD method is only second-order accurate. While second-order convergence may not sound too impressive, we note that in shocked regions we cannot expect more than first-order convergence in any case [92]. Thus, our method is adequate when used in elements with shocks. In this way we obtain benefits of both FV/FD and DG as suggested in [115, 87, 65].
@@
We run a battery of tests with our Nmesh program to verify how well our new method works in practice. We find that the compact FV/FD method indeed reaches second-order accuracy for tests with smooth scalar fields, and that it also performs well in a shock tube test. We then move on to tests with neutron stars in full GRHD. As already described in [76], neutron stars are more challenging due to the presence of vacuum or low-density regions outside the star. We describe in detail the positivity limiters and the atmosphere treatment that we have developed to deal with this challenge. Our combination of limiters and atmosphere are successful in the sense that they enable us to evolve neutron stars with the DG method alone. However, the stars are then plagued by Gibbs oscillations [88, 89, 90, 91] that cause the stars to pulsate. This is especially prominent when the stars are moving. We therefore introduce a troubled-element indicator that uses a list of trouble criteria designed to detect elements that are not giving good results with the DG method. Using this TEI, we build a hybrid scheme that switches between DG and FV/FD as needed. In particular, if an element is marked as troubled and was using DG, we redo the time step with FV/FD. In this way we try to avoid Gibbs oscillations from ever occurring. In case a troubled element is using FV/FD already we remain with FV/FD. On the other hand, if an element is marked as not troubled for 11 consecutive steps and was using FV/FD, we switch to DG, while we remain with DG in case the element was already using it. In DG elements we use only half the number of points per direction as in FV/FD elements, because DG is able to provide higher accuracy results at lower cost due to its higher rate of convergence.
@@
We believe that this is the first time a DG method has been used to evolve the more challenging boosted star (B) and migration (U) cases in full 3D. In case (U), an overdense unstable star gets perturbed to rapidly expand for a while, but later recontracts. The ensuing strong star pulsations contain actual shocks, besides the nonsmooth star surface. Again our hybrid method seems capable of selecting the correct method. The bulk of the star interior and regions further from the star still use the DG method as desired, while FV/FD is selected near the surface and in shocked regions.
@@ Acknowledgements.
It is a pleasure to thank Nils Deppe for helpful discussions. This work was supported by NSF Grant No. PHY-2136036 and No. PHY-2408903 and PeX-FCT (Portugal) Grant No. 2023.12549.PEX. We also acknowledge usage of computer time on the HPC clusters KoKo and Athene at Florida Atlantic University, as well as on Bridges-2, Expanse and Anvil under ACCESS allocations PHY220018 and PHY230109. The figures in this article were created using matplotlib [131, 132] and TikZ [133].
Data Availability
The data that support the findings of this article are openly available [119].
@@ Appendix A Important parameter specifications
Here we state the values of the most important parameters for the different classes of simulations for (S), (P), (U) and (B) that we have discussed in Sec. IV. These are listed in Table 3. By classes, we mean the different sections described above. Hence, they are simply referred to by their section numbers as: §IV.3.1: (S) with pure DG; §IV.3.2: (S) with DG+FV/FD, §IV.4: (P) with pure DG; §IV.5.1: (U) with pure FV/FD and varying p-refinement; §IV.5.2: (U) with pure FV/FD and varying h-refinement; §IV.5.3: (U) with DG+FV/FD; and §IV.6: (B) with DG+FV/FD. We are using the shorthand notation “§” in place of “Sec.” here to denote the sections. Of course Nmesh  has many more parameters that we choose not to list here. When same parameter specification is used across multiple classes, we group those classes together in the table. In cases where the same specification is used for all simulations, we simply say “all” in the “classes” column. Some parameter choices are used for some runs, which are not used for others entirely. In such cases, it is likely that some other parameter is playing an analogous role. For example, when we set the parameter GHG_gammas_type to constants, then we are using the constant constraint damping, and we specify the values of the parameters GHG_gamma0 =0.1=0.1, GHG_gamma1 =1=-1 and GHG_gamma2 =1=1. However, when we set GHG_gammas_type to 3gaussians, as in we set the position dependent triple Gaußian constraint damping functions from Eq. (5), then we have to use the counterpart parameters for these functions. We set these values to:
@@
$A^{(0)}_{(2)}$ and $A^{(2)}_{(2)}$ are set to 0 in our case, as unlike in [97], we have only one star in our simulations. We name the parameters corresponding to the constants $C_{(0)}$, $A^{(0)}_{(1)}$ and $w^{(0)}_{(1)}$ as GHG_3gauss_gamma0_C, GHG_3gauss_gamma0_A1 and GHG_3gauss_gamma0_w1 respectively, and the name of the rest of the parameters follow suite. Owing to the large number of constants, we do not repeat these position-dependent constarint damping parameters in the table. All the run classes that set GHG_gammas_type = 3gaussians use these values for the constants.
@@
by setting amplitude of damped_wave to 0.
@@
GRHD_atmo_cut: Factor $f_{\mathrm{atmo,cut}}$ $\mathtt{1.01\times 10^{-12}}$ §IV.3.1, §IV.4,
from Sec. III.5.1. $\mathtt{10^{-9}}$ §IV.5.1, §IV.3.2, §IV.5.2, §IV.5.3, §IV.6
@@
GRHD_atmo_level: Factor $f_{\mathrm{atmo,level}}$ $\mathtt{10^{-12}}$ §IV.3.1, §IV.4,
from Sec. III.5.1. $\mathtt{10^{-11}}$ §IV.5.1, §IV.3.2, §IV.5.2, §IV.5.3, §IV.6
@@
GRHD_troublescore_cut: Factor $f_{\mathrm{cut,low}}$ related to Eq. (50). $\mathtt{10^{-4}}$ §IV.3.2, §IV.5.3, §IV.6